# Do quantum linear solvers offer advantage for networks-based system of linear equations?

### v3

@@ IV.2 Graph families for incidence matrix-based systems
In this sub-section, we turn our attention to the system of linear equations associated with the incidence matrix of directed graphs. We recall that a sink refers to a vertex whose out-degree is zero and a source to a vertex whose in-degree is zero. Sources and sinks in directed graphs may not be allowed depending on the application of interest. Therefore, for those considered graph families that admit sources and sinks, we consider them as well as their modified versions that do not admit sources and sinks. Our algorithm to convert source and/or sink vertices to non-source and/or sink vertices is presented in Algorithm A.5.2.2 of the Appendix. Table 3 presents the data for all directed graphs considered in this study. We identified 1 best (the directed hypercube graph), 9 better, 4 good, and 6 bad graph families.
IV.2.1 Best graph families
From Table 3, only the directed hypercube graph falls under the best category. Fig. 4 presents the relevant data. As with the hypercube graph family in the Laplacian matrix- based systems, $\kappa$ and $s$ grow as $\mathrm{polylog}(\mathcal{N})$ and $\mathcal{N}\sim n2^{n}$. While the number of vertices still grow as $2^{n}$, the sum of the number of vertices and edges, that is, $\mathcal{N}$, grows as $2^{n}+n2^{n-1}\sim n2^{n}$. Fig. 4(b) shows that phase randomisation, CKS(1), and AQC(1) algorithms begin above the $R(N^{\prime})=1$ line, while HHL approaches this threshold only towards the very end of our $\mathcal{N}$ range.
IV.2.2 Better graph families
We find 9 better graph families in our analysis. We present 3 of them in Fig. 5 and rest of the better graphs can be found in A.11. We observe that similar to the Laplacian matrix case, $\kappa$ and sparsity only grow at most as polylog of $\mathcal{N}$, and the system size growth is sub-exponential but super-linear. Also, just as in the case of Laplacian matrices, CKS(1), AQC(1), and the dream QLS do not graduate any graph families above the better graph family category. In all the better graphs, phase randomisation, CKS(1), and AQC(1) algorithms begin above the $R(\mathcal{N})=1$ line.
IV.2.3 Good graph families
Similar to the system of linear equations associated with the Laplacian matrices, the graph families whose $\kappa$ and/or sparsity grows at most as polylog of $\mathcal{N}$ but whose system size grows linearly only offer a sub-linear speed-up. Fig. A.12 of the Appendix provides the relevant data.
IV.2.4 Bad graph families
Fig. A.13 of the Appendix provides the data for the bad graph families for the system of linear equations associated with the incidence matrices. Either $\kappa$ or $s$ grow polynomially (and in particular, super-linearly), while the system size grows at least linearly. We observe that the dream QLS converts all of the considered graph families from our survey to at least offering a sub-linear advantage, and even offers a polynomial advantage for the GNC graph family.
@@ IV.3 A big-picture view
  • Our numerical survey shows that exponential advantage is uncommon, and polynomial advantage is relatively more common.
  • For all of the best and better graphs that we have identified, the CKS(1) and the AQC(1) algorithms show crossover into the regime of advantage even within the relatively small system sizes that we consider, whereas HHL crosses over only much later or not at all.
@@
  • Our results reiterate the fact that exponential speedup can be witnessed only in classically hard problems. The graph families, which entered the class of best graphs, are classically hard to solve because of their exponential growth in the system size. Graph families in better and good categories have ideal conditions on growth of $\kappa$ and $s$, but their slow growth in system size limits them from achieving exponential speedup with a QLS.
@@ V Combining graph families: generalized hypercube graphs
Our survey underscores the need for finding more graph families with slowly growing $\kappa$ and $s$, but rapidly increasing system size. Motivated by this requirement, we generalize two of the graph families we considered for the Laplacian case to construct the generalized hypercube superfamily.
The idea of hypercubes can be generalized in multiple ways [ghcg]. In this work, we define the vertex set of a generalized hypercube $G_{a}^{m}$ as $V(G_{a}^{m})=\{1,2,3,\cdots,a\}^{\times m}$ which contains $N=a^{m}$ vertices. There is an edge between vertices $x=(x_{1},x_{2},\cdots,x_{m})$ and $y=(y_{1},y_{2},\cdots,y_{m})$ if the Hamming distance $\mathfrak{H}(x,y)=1$. Note that, $x_{i}$ and $y_{i}$ are any of the numbers $1,2,\cdots,a$, for $i=1,2,\cdots,m$. We consider $m$ and $a$ as the parameters of generalized hypercube graphs. The generalized hypercube graphs become the hypercube graphs when $a=2$. These graphs are important as their growth with respect to the parameter $m$ is faster than that of the hypercube graphs. Fig. 6 shows a tableau where we keep $m$ and $a$ on the Y- and X-axes, respectively. Each row or the column in the tableau is a graph family. We term the entire tableau of infinite cells as a generalized hypercube superfamily. The figure has each cell populated with two numbers, one denoting $\kappa$ and the other $s$, which we arrive at by computing the quantities.
@@
    $s$ grows as amm+1am-m+1 (polynomial in $a$).
@@
Although the complete graph is a bad graph, this is not the case for graph families in other rows of Fig. 6. The runtime complexity ratio, $\tilde{R}(a)$ (the variable $a$ plays the role of $n$ from our earlier sections), is $\sim a^{m}\frac{\mathrm{log}(\mathrm{log}(a))}{\mathrm{log}^{2}(a)}$, and in fact grows polynomially.
@@
There exist infinite families of better graphs in the generalized hypercube superfamily. Each row of the tableau, except the first, corresponds to one such family. $\square$
@@
The runtime complexity ratio, $\tilde{R}(m)$ (with $m$ playing the role of $n$ here), is $a^{m}\frac{\mathrm{log}(m)}{m^{11/2}}$, which is exponential.
@@
There exist infinite families of best graphs in the generalized hypercube superfamily, with each column of the tableau corresponding to one such family. $\square$
Having examined the rows and columns of the generalized hypercube superfamily, we now turn to the diagonals of Fig. 6. We discard the first row (which corresponds to complete graph family, which is a bad graph family). On the main diagonal, where the pairs $(a,m)$ correspond to $(a,a)$, the system size scales as $N\sim a^{a}$, while $\kappa=a$ and $s\sim a^{2}$. Thus, $t_{\mathrm{HHL}}\sim a^{7}\mathrm{log}^{2}(a)$ and $t_{\mathrm{CLS}}\sim a^{a}\mathrm{log}(a\mathrm{log}(a))$, and therefore $\tilde{R}(a)$ becomes $\sim a^{a}\mathrm{log}(a\mathrm{log}(a))/\mathrm{log}^{2}(a)$. The main diagonal, thus, constitutes a best graph family.
We next consider super-diagonals and sub-diagonals, that is, diagonals above and below the principal diagonal respectively. The first super-diagonal corresponds to the pair (a,a1)(a,a-1), while the first sub-diagonal corresponds to $(a,a+1)$. In general, cells belonging to the $\mathcal{D}^{th}$ super-diagonal are specified by $(a,a-\mathcal{D})$ and cells belonging to $\mathcal{D}^{th}$ sub-diagonal are represented by $(a,a+\mathcal{D})$. We only consider those super-diagonals and sub-diagonals for which $\mathcal{D}\leq a$.
Claim 3.
For $\mathcal{D}\leq a$, each super-diagonal of the generalized hypercube graph tableau forms a best graph.
Since there is no restriction on the upper limit for $a$ itself, there exist an infinity of such best graph families.
The same ideas apply to the sub-diagonals of the tableau, leading us to our next claim.
Claim 4.
For $\mathcal{D}\leq a$, each sub-diagonal of the generalized hypercube graph tableau forms a best graph.
We conclude by noting that new graph families can be constructed from the generalized hypercube graph tableau by selecting cells that form specific patterns to obtain different degrees of advantage. For example, we can consider iso-$s$ curves. That is, for s=amm+1s=am-m+1, we need to pick $a$ and $m$ such that $s$ is constant. Then, the pairs $(a,m)$ satisfying the condition for different values of $s$ each form a new iso-$s$ bad graph family. We, of course, exclude the solutions corresponding to $s=1$, which are $a=1$ for any $m$ and $m=0$ for any $a$, as they are trivial. For $s>1$, we would be considering solutions of the form $a=\mathfrak{a}+1$ and $m=\frac{s-1}{\mathfrak{a}}$, for $\mathfrak{a}$ being any divisor of s1s-1. Further, the runtime complexities to be a positive quantity, we require $\mathfrak{a}/\mathrm{log}(\mathfrak{a}+1)<s-1$, thereby restricting the number of graphs considered in a iso-$s$ family.
VI Assessing advantage from graphs
@@
    We then evaluate $\kappa$ and $s$ for the given system and study how these quantities scale with system size, besides studying the growth of the system itself.
@@
    Finally, we calculate the runtime ratio, which depends on the interplay between system size growth, $\kappa$, and $s$, and classify the graph family as best, better, good, or bad.
A natural follow-up question is whether one can bypass step 2 and instead qualitatively assess the prospects of an advantage directly from the construction of graph family. In this section, we guess the growth of $\kappa$, $s$ and $\mathcal{N}$, guided by the insights from graph properties, as opposed to quantitatively calculating $\kappa$, which is costly.
For the purposes of this study, we only consider analyses of Laplacian matrix based graph families. We recall that for a graph, the number of vertices $N$ decides the system size for Laplacian matrix $\mathcal{N}$, and $N$ is a function of $n\in\{1,2,3,\cdots\}$. The runtime gain/loss is governed by the growth of $N$, $\kappa$ and $s$ of the Laplacian matrix belonging to a graph family. Now, we elaborate on the properties of a graph family that determine the growth of these parameters which in turn influence the prospects of quantum advantage:
  • System size: The rules of a graph family construction allow one to infer system size growth. For example, in the case of a Sudoku graph (refer A.5.1.1 for its construction), $N$ is restricted to grow as $n^{4}$ by the virtue of its construction. From our observations, we see that degree of runtime advantage is dependent on how fast the system size grows. We observe from Table 2, that a graph family which shows advantage has to grow at least polynomially.
  • Sparsity $s$: The scaling of $s$ can determined by the growth of maximum degree, $d_{max}$, defined by the maximum number of edges incident on any vertex of the graph. The quantity $d_{max}$ is related to $s$ of a Lapalcian matrix as $s=d_{max}+1$. The growth of $s$ can be estimated if one has access to growth of $d_{max}$ either through the definition of graph family or based on numerical analysis. For example, in a complete graph family (see A.5.1.1), the regularity of the graph grows linearly with increase in $N$, which is reflected in the growth of $s$ as well. The data in Table 2 indicates that to realize runtime advantage, $s$ of a graph family must grow sufficiently slowly in $N$.
@@
      The non-zero off-diagonal matrix elements either occur as a scattered distribution (e.g., Figure A.14(b)) or as a number of banded structures (for example, Figure A.14(a)), with increase in system size. We term these kind of matrices as ‘diffuse’. We pick a representative system size and present binary maps of graph families that occur as a diffuse pattern in their adjacency matrix structure in Figures A.14.
@@
      If the number of banded structures remain fixed with the increase in NN (for example, Figure A.15(c) retains 3 bands with increase in $N$), then we term the pattern as ‘sharp’. The binary maps of graph families with sharp structures are presented in Figures A.9.
@@
Given a graph family defined by a construction indexed by $n\in\mathbb{Z}^{+}$, one may expect the family to exhibit a quantum advantage with a QLS for the Laplacian-based NLSP if the following conditions are simultaneously satisfied:
  • the construction leads to a system size growth that is at least polynomial in $n$,
@@
For illustration, we consider the example of the complete graph (see Figure A.14(p)). The details of its construction can be found in Appendix A.5.1.1. The construction leads to the system size scaling being linear, that is, $N=n$. At any system size, each vertex is connected to all other vertices. Thus, $d_{max}$ and hence $s$ grow linearly in $n$. However, since new edges keep spawning from every one of the old vertices for successively increasing system sizes, the family represents the most extreme instance of a diffuse pattern, and thus we also observe a constant value of $\kappa$. Therefore, with a linear growth in system size and linear $s$ scaling, although $\kappa$ is a constant as inferred from the extreme diffuse pattern, it is unlikely that one may realize an advantage with complete graph. Our numerical data in Table 2 shows that it is indeed a bad graph.
On the other hand, when we examine graph families whose constructions involve attaching repeated structural units, such as grid 2d, hexagonal lattice, triangular lattice, ladder, circular ladder, and ring of cliques, and balanced binary tree families, we find the opposite trend. In these cases, since a new structural unit is connected to a constant number of old vertices with the increase in system size, $d_{max}$, and hence $s$ is constant. For the same reason, edge-spawning is not allowed in its construction and thus, $\kappa$ is expected to grow fast according to our conjecture. Therefore, in view of fast $\kappa$ growth in all the graphs, in spite of constant $s$ scaling, achieving advantage is unlikely. This is indeed the case as our data in Table 2 indicates.
@@ VII Quantum hardware demonstrations
Even if a best or a better graph family were available towards some application, realizing a practical quantum advantage requires addressing challenges over and beyond graph theoretic considerations. We list these challenges in Section A.6. Earlier works have already identified the fine print (for examples, see Refs. [morales] and [Aaronson]), and Section A.6 is only added for completeness. In this section, we discuss our quantum hardware results, where we execute toy networks with classically optimized quantum circuits. We also identify the conditions on matrix structure that results in a 1-qubit HHL circuit. These analyses are restricted to Laplacian matrix-based graphs, as incidence matrix-based calculations are not presently feasible on current-day quantum hardware.
@@ VII.0.1 Effective resistance calculation with $4\times 4\ A$ matrices
In this sub-section, we describe our quantum hardware computations on toy networks whose $A$ matrix size is $(4\times 4)$, where we compute effective resistance of electrical circuits shown in Fig. 7 using the HHL algorithm as a representative example, on IonQ Forte-1 quantum computer (one of the current best commercially available computers). We use the HHL algorithm since it is the simplest and most established QLS. The effective resistance is extracted as an overlap between the HHL output vector and a vector with two non-zero entries, 1 and -1, at the vertices of interest. We set $n_{r}=3$ throughout. Due to current-day hardware limitations in terms of their gate fidelities, the use of significant resource reduction methods to reduce circuit depth is required. We borrow resource reduction strategies from Ref. [nishanth2023adapt] for this purpose, and when further resource reduction is necessary, we apply reinforcement learning (RL)-based ZX calculus routine and a causal flow-preserving ZX calculus module discussed and used in Refs. [riu2024reinforcement] and [Palak2025relvqe]. We call this module as RLZX. Hence, our optimization strategy includes the following steps in that sequence: (i) Multi-qubit fixing (introduced in Ref. [nishanth2023adapt]), (ii) Qiskit (version 0.39.5) [Qiskit2021] optimization level 3, which we term L3, (iii) Pytket [Sivarajah2021tket], (iv) Qiskit L3 (version 0.39.5), (v) Qiskit L3 (version 2.0.1), (vi) Approximating unitaries using Qiskit (version 2.0.1), (vii) RLZX, and (viii) Qiskit L3 (version 0.39.5).
In addition to these optimization routines, we also implement the debiasing error mitigation strategy [maksymov2023symmetry] to improve our results. We could only go up to graphs containing $4$ vertices resulting in a $(4\times 4)$ Laplacian matrix, as the resulting HHL quantum circuit for larger problem sizes was too deep to accommodate on the IonQ Forte-1 NISQ era quantum computer even after aggressive resource reduction. All of our circuits are compiled in the $\{RX,\ RY,\ RZ,\ RZZ\}$ gate set. Furthermore, even among the $(4\times 4)$ matrices, we only consider those for which the gate counts post-resource reduction was sufficient for obtaining reasonable results. Table 4 presents the quantum hardware settings as well as the final quality of our result for effective resistance as percentage fraction difference (PFD), that is, our result relative to a classical calculation.

### v4

@@ IV.2 Graph families for incidence matrix-based systems
In this sub-section, we turn our attention to the system of linear equations associated with the incidence matrix of directed graphs. We recall that a sink refers to a vertex whose out-degree is zero and a source to a vertex whose in-degree is zero. Sources and sinks in directed graphs may not be allowed depending on the application of interest. Therefore, for those considered graph families that admit sources and sinks, we consider them as well as their modified versions that do not admit sources and sinks. Our algorithm to convert source and/or sink vertices to non-source and/or sink vertices is presented in Algorithm 1 of the Supplemental Material. Tables 4 and 5 present the data for all directed graphs considered in this study. We identified 14 good and 6 bad graph families from the survey.
IV.2.1 Good graph families
From Table 4, we observe that the growth of $\kappa$ and $s$ is polylogarithmic in $N^{\prime}=N+M$, which is in agreement with the observations from good graph families of Laplacian matrix-based systems. Figs. 4(a) - 4(d) illustrate 2 of the good graphs of this system. The results for rest of the good graphs can be found in Fig. S.10 of the Supplemental Material. Along with the HHL algorithm, these graphs also show exponential advantage with CKS(1), AQC(1), and the dream QLS. Similar to the Laplacian-matrix based systems, the phase randomisation method, CKS(1) and AQC(1) approaches start showing advantage at smaller system sizes compared to the HHL algorithm. We also note that among the good graph families that we identify, the admissible $N^{\prime}$ values in the directed hypercube graph family grow as $N^{\prime}\sim 2^{n}$, and thus the problem is classically hard.
IV.2.2 Bad graph families
Table 5 lists all the graph families that offer no advantage with the HHL algorithm. We also present two instances of the bad graph family in Figs. 4(e)-(h). The rest of the bad graph families can be found in Fig. S.11 of Supplemental Material. Since either $\kappa$ or $s$ or both grow polynomially, we observe that CLS fares well over the HHL algorithm. However, with CKS(1) and AQC(1) methods, 3 bad graph families offer exponential advantage, while with the dream QLS, 4 graph families show exponential advantage.
@@ IV.3 A big-picture view
  • Of the 30 graph families considered for the Laplacian matrix case, 7 show advantage, whereas for the incidence matrix case, 14 out of 20 considered graphs show an advantage.
  • For all of the good graphs that we have identified, the CKS(1) and the AQC(1) algorithms show crossover into the regime of advantage even within the relatively small system sizes that we consider, whereas HHL crosses over only much later or not at all.
@@ V Combining graph families: generalized hypercube graphs
Our survey underscores the need for finding more graph families with slowly growing $\kappa$ and $s$. Motivated by this requirement, we generalize two of the graph families we considered for the Laplacian case to construct the generalized hypercube superfamily.
The idea of hypercubes can be generalized in multiple ways [39]. In this work, we define the vertex set of a generalized hypercube $G_{a}^{m}$ as $V(G_{a}^{m})=\{1,2,3,\cdots,a\}^{\times m}$ which contains $N=a^{m}$ vertices. There is an edge between vertices $x=(x_{1},x_{2},\cdots,x_{m})$ and $y=(y_{1},y_{2},\cdots,y_{m})$ if the Hamming distance $\mathfrak{H}(x,y)=1$. Note that, $x_{i}$ and $y_{i}$ are any of the numbers $1,2,\cdots,a$, for $i=1,2,\cdots,m$. We consider $m$ and $a$ as the parameters of generalized hypercube graphs. The generalized hypercube graphs become the hypercube graphs when $a=2$. These graphs are important as their growth with respect to the parameter $m$ is faster than that of the hypercube graphs. Fig. 5 shows a tableau where we keep $m$ and $a$ on the Y- and X-axes, respectively. Each row or the column in the tableau is a graph family. We term the entire tableau of infinite cells as a generalized hypercube superfamily. The figure has each cell populated with two numbers, one denoting $\kappa$ and the other $s$, which we arrive at by computing the quantities.
@@
    $s$ grows as amm+1am-m+1 (polynomial in $a$). In terms of $N$, $s$ grows as $N^{1/m}$.
@@
Although the complete graph is a bad graph, this is not the case for graph families in other rows of Fig. 5. The runtime complexity ratio for HHL is, ${R}(N)\sim\frac{N^{\alpha}\mathrm{log}(\mathrm{log}(N))}{\mathrm{log(}N)}$ where $0<\alpha<1$. This ensures that there is an exponential separation between the runtimes of HHL algorithm and CLS for every row, which corresponds to a unique graph family of the tableau.
@@
There exist infinite families of good graphs in the generalized hypercube superfamily. Each row of the tableau, except the first, corresponds to one such family. $\square$
@@
The runtime complexity ratio with HHL is, $R(N)\sim N\frac{\mathrm{log}(\log(N))}{\log^{11/2}(N)}$, which is exponential.
@@
There exist infinite families of good graphs in the generalized hypercube superfamily, with each column of the tableau corresponding to one such family. $\square$
VI Assessing advantage from graph family constructions
@@
    We then evaluate $\kappa$ and $s$ for the given system and study how these quantities scale with system size $\mathcal{N}$, which is the number of vertices, $N$, for Laplacian matrix-based systems, and the sum of the number of vertices and edges, $N^{\prime}$, for incidence matrix-based systems. Calculating $\kappa$ is computationally not easy, as it involves finding the largest and smallest eigenvalues of a matrix.
@@
    Finally, we calculate the runtime ratio, which depends on the growth of $\kappa$ and $s$ with $\mathcal{N}$, and classify the graph family as good or bad.
Our survey on 50 graph families led to us finding 21 good graph families, where we could realize an (exponential) advantage with HHL, and 29 bad graph families, where HHL offers no advantage.
A natural follow-up question is whether one can bypass step 2, where computing $\kappa$ for larger system sizes is costly, and instead qualitatively assess the prospects of an advantage with HHL directly from the construction of graph family. We reiterate that analytical expressions for $\kappa$ growth is rare; of the 50 graph families considered, we know $\kappa$ behaviour only for the hypercube graph and the complete graph. In this section, we guess the growth of $\kappa$ and $s$, guided by the insights from graph properties, as opposed to quantitatively calculating $\kappa$.
For the purposes of addressing this question, we only consider analyses of Laplacian matrix based graph families. Thus, the runtime gain/loss is governed by the growth of $\kappa$ and $s$ of the Laplacian matrix belonging to a graph family. Now, we elaborate on the properties of a graph family that determine the growth of these parameters which in turn influence the prospects of quantum advantage:
(a) (b) (c)
(d) (e) (f)
Figure 6: Sub-figures (a), (b) and (c) present the adjacency matrices of representative graphs: hypercube graph, modified Margulis-Gabber-Galil graph and Sudoku graph respectively. These graph families exhibit ‘diffuse’ structure in their adjacency matrix elements. Similarly, sub-figures (d), (e) and (f) that show the matrix element map for grid 2d graph, hexagonal lattice and triangular lattice graphs respectively exhibit ‘sharp’ structure.
  • Sparsity $s$: The scaling of $s$ can be determined by the growth of maximum degree, $d_{max}$, defined by the maximum number of edges incident on any vertex of the graph. The quantity $d_{max}$ is related to $s$ of a Lapalcian matrix as $s=d_{max}+1$. The growth of $s$ can be estimated if one has access to growth of $d_{max}$ either through the definition of graph family or based on numerical analysis. For example, in a complete graph family (see S.5 A 1 of the Supplemental Material), the regularity of the graph grows linearly with increase in $N$, which is reflected in the growth of $s$ as well. As the data in Table 2 indicates, to realize runtime advantage with HHL algorithm, $s$ of a graph family must grow sufficiently slowly in $N$.
@@
      The non-zero off-diagonal matrix elements either occur as a scattered distribution (e.g., 6(b)) or as a number of banded structures (for example, Fig. 6(a) and 6(c)), with increase in system size. We term these kind of matrices as ‘diffuse’. We pick a representative system size and present binary maps of graph families that occur as a diffuse pattern in their adjacency matrix structure for rest of the graph families in Fig. S.12 of Supplemental Material.
@@
      If the number of banded structures remain fixed with the increase in $N$ (for example, Fig. 6(f) retains 3 bands with increase in $N$), then we term the pattern as ‘sharp’. The binary maps for the rest of the graph families with sharp structures are presented in Fig. S.13 of Supplemental Material.
@@
Given a graph family defined by a construction, one may expect the family to exhibit a quantum advantage with HHL for the Laplacian-based NLSP if the following conditions are simultaneously satisfied:
@@
For illustration, we consider the example of the complete graph (see Fig. S.12(m) of Supplemental Material). The details of its construction can be found in Section S.5 A 1. At any system size, $N$, each vertex is connected to all other vertices. Thus, $d_{max}$ and hence $s$ grow linearly in $N$. However, since new edges keep spawning from every one of the old vertices for successively increasing system sizes, the family represents the most extreme instance of a diffuse pattern, and thus we also observe a constant value of $\kappa$. Therefore, although $\kappa$ is a constant as inferred from the extreme diffuse pattern, it is unlikely that one may realize an advantage with HHL for a complete graph given $s$ scales linearly in $N$. Our numerical data in Table 3 shows that it is indeed a bad graph.
On the other hand, when we examine graph families whose constructions involve attaching repeated structural units, such as grid 2d, hexagonal lattice, triangular lattice (plots of adjacency matrices can be found in 6), ladder, circular ladder, and ring of cliques, and balanced binary tree families (plots of adjacency matrices can be found in Fig. S.13 of Supplemental Material), we find the opposite trend. In these cases, since a new structural unit is connected to a constant number of old vertices with the increase in number of vertices, $d_{max}$, and hence $s$ is constant. For the same reason, edge-spawning is not allowed in its construction and thus, $\kappa$ is expected to grow fast according to our conjecture. Therefore, in view of fast $\kappa$ growth in all the graphs, in spite of constant $s$ scaling, achieving advantage with HHL is unlikely. This is indeed the case as our data in Table 3 indicates.
@@ VII Quantum hardware demonstrations
In this section, we discuss our quantum hardware results, where we execute toy networks with classically optimized quantum circuits. We also identify the conditions on matrix structure that results in a 1-qubit HHL circuit. These analyses are restricted to Laplacian matrix-based graphs, as incidence matrix-based calculations are not presently feasible on current-day quantum hardware.
@@ VII.0.1 Effective resistance calculation with $4\times 4\ A$ matrices
In this sub-section, we describe our quantum hardware computations on toy networks whose $A$ matrix size is $(4\times 4)$, where we compute effective resistance of electrical circuits shown in Fig. 7 using the HHL algorithm as a representative example, on IonQ Forte-1 quantum computer (one of the current best commercially available computers). We use the HHL algorithm since it is the simplest and most established QLS. The effective resistance is extracted as an overlap between the HHL output vector and a vector with two non-zero entries, 1 and -1, at the vertices of interest. We set $n_{r}=3$ throughout. Due to current-day hardware limitations in terms of their gate fidelities, the use of significant resource reduction methods to reduce circuit depth is required. We borrow resource reduction strategies from Ref. [7] for this purpose, and when further resource reduction is necessary, we apply reinforcement learning (RL)-based ZX calculus routine and a causal flow-preserving ZX calculus module discussed and used in Refs. [54] and [14]. We call this module as RLZX. Hence, our optimization strategy includes the following steps in that sequence: (i) Multi-qubit fixing (introduced in Ref. [7]), (ii) Qiskit (version 0.39.5) [63] optimization level 3, which we term L3, (iii) Pytket [59], (iv) Qiskit L3 (version 0.39.5), (v) Qiskit L3 (version 2.0.1), (vi) Approximating unitaries using Qiskit (version 2.0.1), (vii) RLZX, and (viii) Qiskit L3 (version 0.39.5).
In addition to these optimization routines, we also implement the debiasing error mitigation strategy [44] to improve our results. We could only go up to graphs containing $4$ vertices resulting in a $(4\times 4)$ Laplacian matrix, as the resulting HHL quantum circuit for larger problem sizes was too deep to accommodate on the IonQ Forte-1 NISQ era quantum computer even after aggressive resource reduction. All of our circuits are compiled in the $\{RX,\ RY,\ RZ,\ RZZ\}$ gate set. Furthermore, even among the $(4\times 4)$ matrices, we only consider those for which the gate counts post-resource reduction was sufficient for obtaining reasonable results. Table 6 presents the quantum hardware settings as well as the final quality of our result for effective resistance as percentage fraction difference (PFD), that is, our result relative to a classical calculation.